# Multispecies inhomogeneous $t$-PushTASEP from antisymmetric fusion

### v2

@@ 6.2. Matrix product formula for the stationary probability
As remarked in the previous subsection, our Theorem 3 reduces the problem of finding the stationary probability of the inhomogeneous $n$-species $t$-PushTASEP to that for a discrete time Markov process whose time evolution is governed by the (suitably normalized) transfer matrix $T^{1}(z|x_{1},\ldots,x_{L})$. Here, we present a simple derivation of the matrix product formula for the stationary probability based on $T^{1}(z|x_{1},\ldots,x_{L})$.
Matrix product formulas were first obtained for homogeneous $n$-species ASEP in [PEM09] using operators defined by nested recursion relations. An inhomogeneous extension was introduced in [CDW15] in connection with the Zamolodchikov-Faddeev algebra and Macdonald polynomials. Further developments on matrix product operators were explored in [KOS24], where the nested recursive structure is identified with the multiline queue construction [CMW22] culminating in a corner transfer matrix formulation of a quantized five-vertex model. It allows for the simplest diagrammatic representation devised to date, with a natural three-dimensional interpretation. [^7] The graphical representation in [CDW15] needs an $n$-color pen, whereas the five-vertex model formulation in [KOS24] uses only two states 00 and $1$. Our presentation here is based on [KOS24].
Let $X_{0}(z),\ldots,X_{n}(z)$ be the “corner transfer matrices” defined in [KOS24, Def.15]. [^8] This is an abuse of terminology from [Bax82, Chap.13], where it is defined for a two-dimensional lattice. Unlike in that context, $X_{i}(z)$ here acts in the direction of a third dimension. These are linear operators depending on the spectral parameter $z$, and act on the $\frac{n(n-1)}{2}$-fold tensor product of $t$-oscillator Fock spaces. To align with the convention used for $R(z)^{\alpha,\beta}_{\gamma,\delta}$ in [KOS24, eq. (16)] and $S^{a,b}_{i,j}(z)$ in (94), we adopt the index transformation $00,1,\ldots,n\rightarrow n,\ldots,1,$ 0.999In this section, we use the simplified notation $S(z)^{a,b}_{i,j}$ for $S(z)^{{\bf e}_{a},{\bf e}_{b}}_{{\bf e}_{i},{\bf e}_{i}}$ as introduced in Appendix A. Further inverting $z$, we set $A_{\alpha}(z)=X_{n-\alpha}(z^{-1})$ for $0\leq\alpha\leq n$. The key result required here is [KOS24, Th.28], which states that the following Zamolodchikov-Faddeev algebra holds:
$\Bigl{(}1-\frac{tz}{x}\Bigr{)}A_{\alpha}(x)A_{\beta}(z)=\sum_{\gamma,\delta=0}^{n}S\Bigl{(}\frac{z}{x}\Bigr{)}^{\beta,\alpha}_{\gamma,\delta}A_{\gamma}(z)A_{\delta}(x).$ (80)
@@
$|\mathbb{P}_{\mathrm{mp}}\rangle=\sum_{(\sigma_{1},\ldots,\sigma_{L})\in\mathcal{S}({\bf m})}{\mathbb{P}}_{\mathrm{mp}}(\sigma_{1},\ldots,\sigma_{L})|\sigma_{1},\ldots,\sigma_{L}\rangle\in\mathbb{V}({\bf m}),$ (81a)
${\mathbb{P}}_{\mathrm{mp}}(\sigma_{1},\ldots,\sigma_{L})=\mathrm{Tr}\left(A_{\sigma_{1}}(x_{1})\cdots A_{\sigma_{L}}(x_{L})\right),$ (81b)
where $\mathbb{V}({\bf m})$ and $\mathcal{S}({\bf m})$ are defined in (5) and (6), respectively. The trace is nonzero and convergent under the assumption $m_{0},\ldots,m_{n}\geq 1$.
Proposition 8.
The vector $|\mathbb{P}_{\mathrm{mp}}\rangle$ is an eigenvector of $T^{1}(z)$ with eigenvalue $\Lambda^{1}(z)$ given by (73). That is,
$T^{1}(z|x_{1},\ldots,x_{L})|\mathbb{P}_{\mathrm{mp}}\rangle=\Lambda^{1}(z|x_{1},\ldots,x_{L})|\mathbb{P}_{\mathrm{mp}}\rangle.$ (82)
@@
From (33), (36b) and (73), the difference between the two sides of (82) is a polynomial in $z$ of degree at most $L$. Therefore it suffices to check the equality at the $L+1$ points $z=0,x_{1},\ldots,x_{L}$. At $z=0$, it follows from (48), (73) and $e_{k-1}(t^{K_{1}},\ldots,t^{K_{n}})+e_{k}(t^{K_{1}},\ldots,t^{K_{n}})=e_{k}(t^{K_{0}},t^{K_{1}},\ldots,t^{K_{n}})$ with $k=1$. (Note $K_{0}=0$.) To verify the equality at the other points, we employ a standard approach. We begin by computing the action of $T^{1}(z|x_{1},\ldots,x_{L})$ using (36b):
$T^{1}(z|x_{1},\ldots,x_{L})|\mathbb{P}_{\mathrm{mp}}\rangle=\sum_{(\sigma_{1},\ldots,\sigma_{L})\in\mathcal{S}({\bf m})}{\mathbb{P}}^{\prime}_{\mathrm{mp}}(\sigma_{1},\ldots,\sigma_{L})|\sigma_{1},\ldots,\sigma_{L}\rangle,$ (83)
@@
${\mathbb{P}}^{\prime}_{\mathrm{mp}}(\sigma_{1},\ldots,\sigma_{L})=\sum_{\begin{subarray}{c}a_{1},\ldots,a_{L}\in\{0,\ldots,n\}\\ (\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L})\in\mathcal{S}({\bf m})\end{subarray}}\!\!\!S\Bigl{(}\frac{z}{x_{1}}\Bigr{)}^{a_{2},\sigma_{1}}_{a_{1},\sigma^{\prime}_{1}}S\Bigl{(}\frac{z}{x_{2}}\Bigr{)}^{a_{3},\sigma_{2}}_{a_{2},\sigma^{\prime}_{2}}\cdots S\Bigl{(}\frac{z}{x_{L}}\Bigr{)}^{a_{1},\sigma_{L}}_{a_{L},\sigma^{\prime}_{L}}\mathrm{Tr}(A_{\sigma^{\prime}_{1}}(x_{1})\cdots A_{\sigma^{\prime}_{L}}(x_{L})).$ (84)
@@
The summations over $(\sigma_{1},\ldots,\sigma_{L})$ and $(\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L})$ are restricted to $\mathcal{S}({\bf m})$ by the weight conservation property of $T^{1}(z)$ and $S(z)$. Now, consider the specialization $z=x_{1}$. From (90), the leftmost factor $S(z/x_{1})^{a_{2},\sigma_{1}}_{a_{1},\sigma^{\prime}_{1}}$ in (84) simplifies to $(1-t)\delta_{a_{1},\sigma_{1}}\delta_{a_{2},\sigma^{\prime}_{1}}$. Substituting this into the RHS of (84) gives
@@
$(1-t)\sum_{\begin{subarray}{c}a_{3},\ldots,a_{L}\\ \sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}\end{subarray}}\!\!\!S\Bigl{(}\frac{x_{1}}{x_{2}}\Bigr{)}^{a_{3},\sigma_{2}}_{\sigma^{\prime}_{1},\sigma^{\prime}_{2}}S\Bigl{(}\frac{x_{1}}{x_{3}}\Bigr{)}^{a_{4},\sigma_{3}}_{a_{3},\sigma^{\prime}_{3}}\cdots S\Bigl{(}\frac{x_{1}}{x_{L}}\Bigr{)}^{\sigma_{1},\sigma_{L}}_{a_{L},\sigma^{\prime}_{L}}\mathrm{Tr}(A_{\sigma^{\prime}_{1}}(x_{1})A_{\sigma^{\prime}_{2}}(x_{2})\cdots A_{\sigma^{\prime}_{L}}(x_{L})).$ (85)
Applying (80), we sum over $\sigma^{\prime}_{1},\sigma^{\prime}_{2}$ obtaining
@@
$(1-t)(1-\frac{tx_{1}}{x_{2}})\sum_{\begin{subarray}{c}a_{3},\ldots,a_{L}\\ \sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}\end{subarray}}\!\!\!S\Bigl{(}\frac{x_{1}}{x_{3}}\Bigr{)}^{a_{4},\sigma_{3}}_{a_{3},\sigma^{\prime}_{3}}\cdots S\Bigl{(}\frac{x_{1}}{x_{L}}\Bigr{)}^{\sigma_{1},\sigma_{L}}_{a_{L},\sigma^{\prime}_{L}}\mathrm{Tr}(A_{\sigma_{2}}(x_{2})A_{a_{3}}(x_{1})A_{\sigma^{\prime}_{3}}(x_{3})\cdots A_{\sigma^{\prime}_{L}}(x_{L})).$ (86)
@@
We can successively push $A_{\bullet}(x_{1})$ with any index $\bullet$ to the right using (80), leading to
@@
$\prod_{j=1}^{L}\left(1-\frac{tx_{1}}{x_{j}}\right)\mathrm{Tr}\left(A_{\sigma_{2}}(x_{2})A_{\sigma_{3}}(x_{3})\cdots A_{\sigma_{L}}(x_{L})A_{\sigma_{1}}(x_{1})\right)=\Lambda^{1}(x_{1}|x_{1},\ldots,x_{L}){\mathbb{P}}_{\mathrm{mp}}(\sigma_{1},\ldots,\sigma_{L}).$ (87)
@@
From Proposition 8, it follows that the matrix product state (81a) is a joint eigenstate of $T^{0}(z),\ldots,T^{n+1}(z)$. Given their eigenvalues $\Lambda^{0}(z),\ldots,\Lambda^{n+1}(z)$ as in (73), along with the result (78) and the uniqueness of the stationary state, we conclude that (81b) provides a matrix product formula for the (unnormalized) stationary probability of the inhomogeneous $n$-species $t$-PushTASEP.
@@ 7. ASEP Markov matrix from transfer matrix
For readers convenience, we include a short elementary section recalling the well-known origin of the $n$-species ASEP Markov matrix in a commuting family of transfer matrices in the convention of this paper. The ASEP is another Markov process on each sector $\mathbb{V}({\bf m})$ in (5). Its Markov matrix consists of the nearest neighbor interaction terms as
$H_{\text{ASEP}}$ $=\sum_{i\in{\mathbb{Z}}_{L}}1\otimes\cdots\otimes 1\otimes h_{\text{ASEP}}\otimes 1\otimes\cdots\otimes 1,$ (88a)
$h_{\text{ASEP}}({\mathsf{v}}_{\alpha}\otimes{\mathsf{v}}_{\beta})$ $=({\mathsf{v}}_{\beta}\otimes{\mathsf{v}}_{\alpha}-{\mathsf{v}}_{\alpha}\otimes{\mathsf{v}}_{\beta})t^{[\alpha>\beta]}.$ (88b)
where in (88a), $h_{\text{ASEP}}$ acts on the $(i,i+1)$ components of ${\mathsf{V}}^{\otimes L}$. It swaps the local states $0\leq\alpha,\beta\leq n$ in adjacent sites with the rate $t^{[\alpha>\beta]}$.
Let us consider the transfer matrix $T^{k}(z)$ in (36a)–(36b) in the special case $k=1$ with the homogeneous choice of parameters $x_{1}=\cdots=x_{L}=1$. We denote it as $T^{1}(z|{\bf x}={\bf 1})$. As a corollary of (37), they still satisfy the commutativity:
$[T^{1}(z|{\bf x}={\bf 1}),T^{1}(z^{\prime}|{\bf x}={\bf 1})]=0.$ (89)
For the simplest $R$-matrix $S(z)=S^{1,1}(z)$ in (94), which is relevant to $T^{1}(z|{\bf x}={\bf 1})$, it is straightforward to check
$S(1)$ $=(1-t)\mathcal{P},\quad\mathcal{P}({\mathsf{v}}\otimes{\mathsf{v}}^{\prime})={\mathsf{v}}^{\prime}\otimes{\mathsf{v}},$ (90)
$\mathcal{P}\!\left.\frac{dS(z)}{dz}\right|_{z=1}$ $=-h_{\text{ASEP}}-t\,\mathrm{Id}.$ (91)
From (90), one finds that $T^{1}(1|{\bf x}={\bf 1})=(1-t)^{L}\mathcal{C}$, where $\mathcal{C}({\mathsf{v}}_{\sigma_{1}}\otimes{\mathsf{v}}_{\sigma_{2}}\otimes\cdots\otimes{\mathsf{v}}_{\sigma_{L}})={\mathsf{v}}_{\sigma_{L}}\otimes{\mathsf{v}}_{\sigma_{1}}\otimes\cdots\otimes{\mathsf{v}}_{\sigma_{L-1}}$ represents a cyclic shift. Using this result, (91) leads, via an argument analogous to [KMMO16, eq. (55)], to
$H_{\text{ASEP}}=-(1-t)\frac{d}{dz}\left.\log T^{1}(z|{\bf x}={\bf 1})\right|_{z=1}-tL\,\mathrm{Id},$ (92)
@@
Recall the joint eigenvector |¯$|\overline{\mathbb{P}}({\bf m})\rangle$ of $T^{0}(z),\ldots,T^{n+1}(z)$ with eigenvalues $\Lambda^{0}(z),\ldots,\Lambda^{n+1}(z)$ introduced in Section 6.1. They all depend on the inhomogeneities $x_{1},\ldots,x_{L}$. From the specialization $x_{1}=\cdots=x_{L}=1$ and (73), we have
$T^{1}(z|{\bf x}={\bf 1})|\overline{\mathbb{P}}({\bf m})\rangle_{{\bf x}={\bf 1}}$ $=\Lambda^{1}(z|{\bf x}={\bf 1})|\overline{\mathbb{P}}({\bf m})\rangle_{{\bf x}={\bf 1}},$ (93a)
$\Lambda^{1}(z|{\bf x}={\bf 1})$ $=(1-tz)^{L}+e_{1}(t^{K_{1}},\ldots,t^{K_{n}})(1-z)^{L}.$ (93b)

### v3

@@ 6.2. Matrix product formula for the stationary probability
As remarked in the previous subsection, our Theorem 5.1 reduces the problem of finding the stationary probability of the inhomogeneous $n$-species $t$-PushTASEP to that for a discrete time Markov process whose time evolution is governed by the (suitably normalized) transfer matrix $T^{1}(z|x_{1},\ldots,x_{L})$. Here, we present a simple derivation of the matrix product formula for the stationary probability based on $T^{1}(z|x_{1},\ldots,x_{L})$.
Matrix product formulas were first obtained for homogeneous $n$-species ASEP in [PEM09] using operators defined by nested recursion relations. An inhomogeneous extension was introduced in [CDW15] in connection with the Zamolodchikov-Faddeev algebra and Macdonald polynomials. Further developments on matrix product operators were explored in [KOS24], where the nested recursive structure is identified with the multiline queue construction [CMW22] culminating in a corner transfer matrix formulation of a quantized five-vertex model. It allows for the simplest diagrammatic representation devised to date, with a natural three-dimensional interpretation. [^6] The graphical representation in [CDW15] needs an $n$-color pen, whereas the five-vertex model formulation in [KOS24] uses only two states 0 and $1$. Our presentation here is based on [KOS24].
Let $X_{0}(z),\ldots,X_{n}(z)$ be the “corner transfer matrices” defined in [KOS24, Def.15]. [^7] This is an abuse of terminology from [Bax82, Chap.13], where it is defined for a two-dimensional lattice. Unlike in that context, $X_{i}(z)$ here acts in the direction of a third dimension. These are linear operators depending on the spectral parameter $z$, and act on the $\frac{n(n-1)}{2}$-fold tensor product of $t$-oscillator Fock spaces. To align with the convention used for $R(z)^{\alpha,\beta}_{\gamma,\delta}$ in [KOS24, eq. (16)] and $S^{a,b}_{i,j}(z)$ in (3.6), we adopt the index transformation $0,1,\ldots,n\rightarrow n,\ldots,1,0$. [^8] In this section, we use the simplified notation $S(z)^{a,b}_{i,j}$ for $S(z)^{{\bf e}_{a},{\bf e}_{b}}_{{\bf e}_{i},{\bf e}_{i}}$ as introduced in Section 3.2. Further inverting $z$, we set $A_{\alpha}(z)=X_{n-\alpha}(z^{-1})$ for $0\leq\alpha\leq n$. The key result required here is [KOS24, Th.28], which states that the following Zamolodchikov-Faddeev algebra holds:
$\Bigl(1-\frac{tz}{x}\Bigr)A_{\alpha}(x)A_{\beta}(z)=\sum_{\gamma,\delta=0}^{n}S\Bigl(\frac{z}{x}\Bigr)^{\beta,\alpha}_{\gamma,\delta}A_{\gamma}(z)A_{\delta}(x).$ (6.7)
@@
$|\mathbb{P}_{\mathrm{mp}}\rangle=\sum_{(\sigma_{1},\ldots,\sigma_{L})\in\mathcal{S}({\bf m})}{\mathbb{P}}_{\mathrm{mp}}(\sigma_{1},\ldots,\sigma_{L})|\sigma_{1},\ldots,\sigma_{L}\rangle\in\mathbb{V}({\bf m}),$ (6.8a)
${\mathbb{P}}_{\mathrm{mp}}(\sigma_{1},\ldots,\sigma_{L})=\mathrm{Tr}\left(A_{\sigma_{1}}(x_{1})\cdots A_{\sigma_{L}}(x_{L})\right),$ (6.8b)
where $\mathbb{V}({\bf m})$ and $\mathcal{S}({\bf m})$ are defined in (2.1) and (2.2), respectively. The trace is nonzero and convergent under the assumption $m_{0},\ldots,m_{n}\geq 1$.
Proposition 6.2.
The vector $|\mathbb{P}_{\mathrm{mp}}\rangle$ is an eigenvector of $T^{1}(z)$ with eigenvalue $\Lambda^{1}(z)$ given by (6.1). That is,
$T^{1}(z|x_{1},\ldots,x_{L})|\mathbb{P}_{\mathrm{mp}}\rangle=\Lambda^{1}(z|x_{1},\ldots,x_{L})|\mathbb{P}_{\mathrm{mp}}\rangle.$ (6.9)
@@
From (3.12), (4.2b) and (6.1), the difference between the two sides of (6.9) is a polynomial in $z$ of degree at most $L$. Therefore it suffices to check the equality at the $L+1$ points $z=0,x_{1},\ldots,x_{L}$. At $z=0$, it follows from (4.14), (6.1) and $e_{k-1}(t^{K_{1}},\ldots,t^{K_{n}})+e_{k}(t^{K_{1}},\ldots,t^{K_{n}})=e_{k}(t^{K_{0}},t^{K_{1}},\ldots,t^{K_{n}})$ with $k=1$. (Note $K_{0}=0$.) To verify the equality at the other points, we employ a standard approach. We begin by computing the action of $T^{1}(z|x_{1},\ldots,x_{L})$ using (4.2b):
$T^{1}(z|x_{1},\ldots,x_{L})|\mathbb{P}_{\mathrm{mp}}\rangle=$ $\sum_{(\sigma_{1},\ldots,\sigma_{L})\in\mathcal{S}({\bf m})}{\mathbb{P}}^{\prime}_{\mathrm{mp}}(\sigma_{1},\ldots,\sigma_{L})|\sigma_{1},\ldots,\sigma_{L}\rangle,$ (6.10)
@@
${\mathbb{P}}^{\prime}_{\mathrm{mp}}(\sigma_{1},\ldots,\sigma_{L})=$ $\sum_{\begin{subarray}{c}a_{1},\ldots,a_{L}\in\{0,\ldots,n\}\\ (\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L})\in\mathcal{S}({\bf m})\end{subarray}}\!\!\!S\Bigl(\frac{z}{x_{1}}\Bigr)^{a_{2},\sigma_{1}}_{a_{1},\sigma^{\prime}_{1}}S\Bigl(\frac{z}{x_{2}}\Bigr)^{a_{3},\sigma_{2}}_{a_{2},\sigma^{\prime}_{2}}\cdots S\Bigl(\frac{z}{x_{L}}\Bigr)^{a_{1},\sigma_{L}}_{a_{L},\sigma^{\prime}_{L}}$ (6.11)
$\times\mathrm{Tr}(A_{\sigma^{\prime}_{1}}(x_{1})\cdots A_{\sigma^{\prime}_{L}}(x_{L})).$ (6.12)
@@
The summations over $(\sigma_{1},\ldots,\sigma_{L})$ and $(\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L})$ are restricted to $\mathcal{S}({\bf m})$ by the weight conservation property of $T^{1}(z)$ and $S(z)$. Now, consider the specialization $z=x_{1}$. From (7.3), the leftmost factor $S(z/x_{1})^{a_{2},\sigma_{1}}_{a_{1},\sigma^{\prime}_{1}}$ in (6.12) simplifies to $(1-t)\delta_{a_{1},\sigma_{1}}\delta_{a_{2},\sigma^{\prime}_{1}}$. Substituting this into the RHS of (6.12) gives
@@
$(1-t)\sum_{\begin{subarray}{c}a_{3},\ldots,a_{L}\\ \sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}\end{subarray}}\!\!\!S\Bigl(\frac{x_{1}}{x_{2}}\Bigr)^{a_{3},\sigma_{2}}_{\sigma^{\prime}_{1},\sigma^{\prime}_{2}}S\Bigl(\frac{x_{1}}{x_{3}}\Bigr)^{a_{4},\sigma_{3}}_{a_{3},\sigma^{\prime}_{3}}\cdots S\Bigl(\frac{x_{1}}{x_{L}}\Bigr)^{\sigma_{1},\sigma_{L}}_{a_{L},\sigma^{\prime}_{L}}\mathrm{Tr}(A_{\sigma^{\prime}_{1}}(x_{1})A_{\sigma^{\prime}_{2}}(x_{2})\cdots A_{\sigma^{\prime}_{L}}(x_{L})).$ (6.13)
Applying (6.7), we sum over $\sigma^{\prime}_{1},\sigma^{\prime}_{2}$ obtaining
@@
$(1-t){\left(1-\frac{tx_{1}}{x_{2}}\right)}\sum_{\begin{subarray}{c}a_{3},\ldots,a_{L}\\ \sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}\end{subarray}}S\Bigl(\frac{x_{1}}{x_{3}}\Bigr)^{a_{4},\sigma_{3}}_{a_{3},\sigma^{\prime}_{3}}\cdots S\Bigl(\frac{x_{1}}{x_{L}}\Bigr)^{\sigma_{1},\sigma_{L}}_{a_{L},\sigma^{\prime}_{L}}\\ \times\mathrm{Tr}(A_{\sigma_{2}}(x_{2})A_{a_{3}}(x_{1})A_{\sigma^{\prime}_{3}}(x_{3})\cdots A_{\sigma^{\prime}_{L}}(x_{L})).$ (6.14)
@@
We can successively push $A_{\bullet}(x_{1})$ with any index $\bullet$ to the right using (6.7), leading to
@@
$\prod_{j=1}^{L}\left(1-\frac{tx_{1}}{x_{j}}\right)\mathrm{Tr}\left(A_{\sigma_{2}}(x_{2})A_{\sigma_{3}}(x_{3})\cdots A_{\sigma_{L}}(x_{L})A_{\sigma_{1}}(x_{1})\right)\\ =\Lambda^{1}(x_{1}|x_{1},\ldots,x_{L}){\mathbb{P}}_{\mathrm{mp}}(\sigma_{1},\ldots,\sigma_{L}).$ (6.15)
@@
From Proposition 6.2, it follows that the matrix product state (6.8a) is a joint eigenstate of $T^{0}(z),\ldots,\allowbreak T^{n+1}(z)$. Given their eigenvalues $\Lambda^{0}(z),\ldots,\Lambda^{n+1}(z)$ as in (6.1), along with the result (6.6) and the uniqueness of the stationary state, we conclude that (6.8b) provides a matrix product formula for the (unnormalized) stationary probability of the inhomogeneous $n$-species $t$-PushTASEP.
@@ 7. ASEP Markov matrix from transfer matrix
For readers convenience, we include a short elementary section recalling the well-known origin of the $n$-species ASEP Markov matrix in a commuting family of transfer matrices in the convention of this paper. The ASEP is another Markov process on each sector $\mathbb{V}({\bf m})$ in (2.1). Its Markov matrix consists of the nearest neighbor interaction terms as
$H_{\text{ASEP}}$ $=\sum_{i\in{\mathbb{Z}}_{L}}1\otimes\cdots\otimes 1\otimes h_{\text{ASEP}}\otimes 1\otimes\cdots\otimes 1,$ (7.1a)
$h_{\text{ASEP}}({\mathsf{v}}_{\alpha}\otimes{\mathsf{v}}_{\beta})$ $=({\mathsf{v}}_{\beta}\otimes{\mathsf{v}}_{\alpha}-{\mathsf{v}}_{\alpha}\otimes{\mathsf{v}}_{\beta})t^{[\alpha>\beta]}.$ (7.1b)
where in (7.1a), $h_{\text{ASEP}}$ acts on the $(i,i+1)$ components of ${\mathsf{V}}^{\otimes L}$. It swaps the local states $0\leq\alpha,\beta\leq n$ in adjacent sites with the rate $t^{[\alpha>\beta]}$.
Let us consider the transfer matrix $T^{k}(z)$ in (4.2a)–(4.2b) in the special case $k=1$ with the homogeneous choice of parameters $x_{1}=\cdots=x_{L}=1$. We denote it as $T^{1}(z|{\bf x}={\bf 1})$. As a corollary of (4.3), they still satisfy the commutativity:
$[T^{1}(z|{\bf x}={\bf 1}),T^{1}(z^{\prime}|{\bf x}={\bf 1})]=0.$ (7.2)
For the simplest $R$-matrix $S(z)=S^{1,1}(z)$ in (3.6), which is relevant to $T^{1}(z|{\bf x}={\bf 1})$, it is straightforward to check
$S(1)$ $=(1-t)\mathcal{P},\quad\mathcal{P}({\mathsf{v}}\otimes{\mathsf{v}}^{\prime})={\mathsf{v}}^{\prime}\otimes{\mathsf{v}},$ (7.3)
$\mathcal{P}\!\left.\frac{dS(z)}{dz}\right|_{z=1}$ $=-h_{\text{ASEP}}-t\,\mathrm{Id}.$ (7.4)
From (7.3), one finds that $T^{1}(1|{\bf x}={\bf 1})=(1-t)^{L}\mathcal{C}$, where $\mathcal{C}({\mathsf{v}}_{\sigma_{1}}\otimes{\mathsf{v}}_{\sigma_{2}}\otimes\cdots\otimes{\mathsf{v}}_{\sigma_{L}})={\mathsf{v}}_{\sigma_{L}}\otimes{\mathsf{v}}_{\sigma_{1}}\otimes\cdots\otimes{\mathsf{v}}_{\sigma_{L-1}}$ represents a cyclic shift. Using this result, (7.4) leads, via an argument analogous to [KMMO16, eq. (55)], to
$H_{\text{ASEP}}=-(1-t)\frac{d}{dz}\left.\log T^{1}(z|{\bf x}={\bf 1})\right|_{z=1}-tL\,\mathrm{Id},$ (7.5)
@@
Recall the joint eigenvector $|\overline{\mathbb{P}}({\bf m})\rangle$ of $T^{0}(z),\ldots,T^{n+1}(z)$ with eigenvalues $\Lambda^{0}(z),\ldots,\allowbreak\Lambda^{n+1}(z)$ introduced in Section 6.1. They all depend on the inhomogeneities $x_{1},\ldots,x_{L}$. From the specialization $x_{1}=\cdots=x_{L}=1$ and (6.1), we have
$T^{1}(z|{\bf x}={\bf 1})|\overline{\mathbb{P}}({\bf m})\rangle_{{\bf x}={\bf 1}}$ $=\Lambda^{1}(z|{\bf x}={\bf 1})|\overline{\mathbb{P}}({\bf m})\rangle_{{\bf x}={\bf 1}},$ (7.6a)
$\Lambda^{1}(z|{\bf x}={\bf 1})$ $=(1-tz)^{L}+e_{1}(t^{K_{1}},\ldots,t^{K_{n}})(1-z)^{L}.$ (7.6b)